OpenStudy (anonymous):

The position of the digits of a two digit number is interchanged. If the sum of the number and the number obtained by interchanging the positions is 55, what is the sum of the digits of the number?

OpenStudy (unklerhaukus):

five

OpenStudy (anonymous):

true and thenx . .Let the unit's digit of the number be x and its ten's digit y.<br>Number = 10y + x, when the digits interchange their positions, new number formed is 10x + y. It is given, 10y + x + 10x + y = 55 or <br>11x + 11y = 55 or 11(x + y) = 55 Or x + Y = 55/11 or x + y = 5. Hence the sum of the digits is 5.

OpenStudy (unklerhaukus):

i just guessed a solution, and checked if it worked, 23+32=55, 14+41=55 also works

OpenStudy (unklerhaukus):

your solution is much more mathematical.